PHY-1103 Fall 2018 Cole KCh7 HW Problem 7.14 9 of 11> rarL A What is the speed of each section (relative to Earth) after the explosion? onstants A735-kg two-stage rocket is traveling at a speed of 6.80x103 m/s away from Earth when a predesigned explosion separates the rocket into two sections of equal mass that then move with a speed of 2.60x10 m/s relative to each other along the original line of motion. Express your answers using three significant figures separated by a comma ΑΣφ m/s Submit Request Answer Part B What are the direction of each section (relative to Earth) after the explosion? both away from Earth both towards the Earth one away from Earth and the other towards the EarthA 735-kg two-stage rocket is traveling at a speed of 6.80×103 m/s away from Earth when a predesigned explosion separates the rocket into two sections of equal mass that then move with a speed of 2.60×103 m/s relative to each other along the original line of motion.
Answer #1

A)

Using law of conservation of momentum,

mv = 0.5 mv1 + 0.5 mv2

2v = v1 + v2

2 x 6800 = 2 v1 - 2600

V1 = 8100 m/s

V2 = 8100 - 2600 = 5500 m/s

(8100, 5500) m/s

B)

In the direction of original line of motion...

C)

E = 0.5 x (735 x 8100^2/2) + 0.5 x (735 x 5500^2/2) - (0.5 x 735 x 6800^2)

E = 6.21 x 10^8 J

Comment in case any doubt please rate my answer...